OpenStudy (anonymous):

Write the definition of the following statement about sets? {n^2 + n + 1: n∈ℕ } ⊆ {2n + 1 : n∈ℕ } the answer provided is: ∀n∈ℕ ∃m∈ℕ (n^2 + n + 1 = 2m + 1} The answer intuitively makes sense to me, but I don't know how it was derived What I tried was x∈{n^2 + n + 1: n∈ℕ } → x∈ {2n + 1 : n∈ℕ } ∃n∈ℕ (x = n^2 + n + 1} → ∃m∈ℕ (x = 2m + 1) But i'm stuck here and can't get it to match the answer provided

OpenStudy (experimentx):

this is essentially saying that every number of the n^2 + n + 1 can be written as 2m+1 or just prove n^2 + n + 1 is always odd.

OpenStudy (experimentx):

take two cases ... n=2i or n=2i+1

OpenStudy (experimentx):

well well that's better.

OpenStudy (anonymous):

\[ n^2 + n+1=n(n+1) +1 \] n(n+1) is always even hence n(n+1) +1 is odd

OpenStudy (anonymous):

well I realized n^2 + n + 1 is always odd. What i'm confused about is sort of like why the answer was the way it is. I.e ∀n∈ℕ ∃m∈ℕ (n^2 + n + 1 = 2m + 1} What steps that leads to the introducing the quantifiers ?

OpenStudy (anonymous):

I prefer to write it out step by step that lead to ∀n∈ℕ ∃m∈ℕ (n^2 + n + 1 = 2m + 1} Noticing that n^2 + n + 1 is odd somehow doesn't provide the connection between the two sets to me.

OpenStudy (experimentx):

Let, \(m = 1 + 2 + + \dots +n =\frac{n(n+1)}{2}\) then \( \forall n \in \Bbb N , \exists m \in \Bbb N | (n^2+n+1 = 2m+1) \)

OpenStudy (anonymous):

huhm... This works fine